OpenStudy (ray1998):

Last question, I promise. What is the solution of the linear quadratic system of equations? y = x^2 + 5x - 3 y - x = 2 I need to be walked through it, not just given the answer. Will give fan and medal. Thank you in advance! :)

OpenStudy (ray1998):

Do you plug the 2 in like this? y = 2^2 + 5(2) - 3

OpenStudy (ybarrap):

Not exactly... check this out: So you have $$ y-x=2\\ y=2+x\\ $$ Plug in to your quadratic $$ y=x^2+5x-3=2+x\\ x^2+5x-x-3-2=0\\ x^2+4x-5=0\\ (x-1)(x+5)=0 $$ So there are two solutions if x=1 the equation is true if x=-5 the equation is true We know that y=2+x So when x=1, y =3 and when x=-5, y = -3 Two solutions: $$ (1,3)\text{ and }(-5,-3) $$ Make sense?

OpenStudy (ray1998):

Sorta. I'll have to look at it some more, but I get part of it. Thank you. :)

OpenStudy (ybarrap):